### v2

@@ Abstract
The Contrastive Language-Image Pretraining (CLIP) model has been widely used in various downstream vision tasks. To augment its capacity for these tasks, the few-shot learning paradigm has been widely adopted. However, current paradigms may struggle with fine-grained classification and satellite image recognition due to domain gaps widening. To address this limitation, we propose retrieval-enhanced visual prompt learning (RePrompt), which introduces retrieval mechanisms to cache the knowledge of downstream tasks. RePrompt constructs a retrieval database from either training examples or external data if available, and uses a retrieval mechanism to enhance multiple stages of a simple prompt learning baseline, thus narrowing the domain gap. During inference, our enhanced model can reference similar samples brought by retrieval to make more accurate predictions. A detailed analysis reveals that retrieval helps to improve the distribution of late features, thus, improving generalization for downstream tasks. Reprompt attains state-of-the-art performance on a wide range of vision datasets, including 11 image datasets, 3 video datasets, 1 multi-view dataset, and 4 domain generalization benchmarks.
Keywords:
@@
few-shot learning prompt tuning image classification video understanding multi-view image classification
1 Introduction
@@
Visual concept recognition has achieved remarkable success in closed-set scenarios with large-scale training sets, typically ImageNet dataset. The recognition accuracy can even surpass human ability. However, it may not always be feasible to have a large training set for each specific visual concept in certain downstream tasks. How to learn a robust visual concept recognition system with low-shot data has become a challenging but valuable problem in computer vision. Previous methods MAML ; prototypical ; matching mainly focus on learning a transferable visual representation from a source domain and quickly adapting to few-shot downstream tasks through fine-tuning techniques. However, limited by the categories and closed-set training samples in the source domain, these few-shot learning algorithms are only effective in impractically simple settings, such as distinguishing $5$-way $1$-shot classification.
@@
Recent advancements in visual representation learning have been propelled by the emergence of large-scale vision-language models, such as CLIP CLIP and ALIGN ALIGN . The feasibility of applying these visual language models in addressing more challenging low-shot learning problems has attracted considerable attention within the research community CLIP ; CoOP ; clip_adapter ; video_prompt . In practice, it is cumbersome to fine-tune the entire vision-language model for transfer learning in each downstream task. Furthermore, finetuning the full model can result in catastrophic forgetting catastrophic_forgetting ; example_forgetting , leading to poor performance on new tasks despite the pre-trained model’s initial success. To address these challenges, CoOP CoOP first proposes a few-shot evaluation framework for visual language models. There are full $C$ classes of downstream tasks and each class has 1/2/4/8/16-shot samples. Meanwhile, CoOP designs a learnable text prompt to replace the sub-optimal hand-crafted text prompt templates.
@@
Nevertheless, the frozen image feature representation also leads to sub-optimal performance. Inspired by VPT VPT , many studies UPT ; MaPLe ; CAVPT ; video_prompt ; promptsrc consider adding additional learnable visual tokens into the image encoder to parameter-efficient fine-tune the vision-language models on downstream tasks. Despite significant improvements achieved by these methods for few-shot learning, these parametric models struggle to generalize to extremely low-shot data or atypical instances, often relying on rote memorization. For instance, given only one image per class, CoOP performs even worse than the zero-shot classification results of CLIP.
@@
Retrieval-augmentation approaches retrieval_long_tail ; RETRO_tokens ; RETROPROMPT ; retrieval_pretraining ; Re-ViLM ; tip-adapter are employed to retrieve knowledge corpus and generate additional references, thereby enhancing performance in low-resource scenarios. Large Language Models (LLM) also get advancements with retrieval-augmentation approaches and further develop Retrieval Augmentation Generation (RAG) RAG . However, the application of retrieval augmentation to visual language models remains underexplored. Tip-Adapter  tip-adapter , a key-value cache model, is constructed as a retrieval auxiliary classifier for CLIP. It leverages the similarity between the test and few-shot training samples, calculating affinity matrixes in text feature space and image feature space. However, it is observed that the variances and means of visual modal similarities are smaller compared to those in visual-language modalities udandarao2022sus . This phenomenon is caused by CLIP maximizing the cosine similarity between paired samples in different modalities during its contrastive learning training, without considering the intra-modality similarity. Consequently, the image similarity metrics employed by Tip-Adapter, derived from CLIP, are potentially compromised by domain gaps and an insufficient adaptation to downstream datasets.
@@
In this work, we investigate the feasibility of learning prompts from a cache model, which facilitates the enhancement of query image features to adjust image-image similarities. To simultaneously optimize these aspects, we introduce Retrieval-enhanced visual Prompt learning (RePrompt), which incorporates a retrieval mechanism for associatively learning knowledge representations from downstream tasks. RePrompt establishes a retrieval database that stores common training image representations as retrieval keys. The current image representations retrieve relevant cached knowledge from the database. Unlike Sus-X udandarao2022sus , which expands the retrieval database using external datasets like LAION schuhmann2022laion , our method adheres strictly to the principles of few-shot learning by not utilizing external datasets. Nonetheless, our method can still benefit from incorporating external knowledge.
@@ 2 Related Work
Retrieval-augmented models have gained significant attention in the fields of Natural Language Processing (NLP) and Computer Vision (CV). Numerous studies have augmented large language models with external memory through retrieval-enhancement mechanism dense_retrieval ; retrieval_pretraining ; bert_knn ; RETRO_tokens ; RETROPROMPT ; retrieval_generation . The integration of knowledge into a language model involves retrieving relevant samples from external memory, thereby enabling the model to generate more informed predictions based on these samples. Typically, the external memory comprises a collection of text paragraphs or a structured knowledge base. The BERT-kNN model bert_knn combines the outputs of a trained BERT model with a non-parametric $k$NN model for question answering tasks. REALM retrieval_pretraining acquires external knowledge from Wikipedia. Karpukhin et al. dense_retrieval developed a dense passage retrieval system for open-domain question answering. RETRO RETRO_tokens demonstrates how to blend retrieved tokens with input tokens in a transformer architecture and methodically evaluates the influence of large-scale external memory datasets on NLP tasks. Chen et al. RETROPROMPT utilize retrieved tokens as input prompts and fine-tune the entire language model. Similar to the approach in RETROPROMPT , we employ retrieval samples as prompts to introduce pertinent information.
@@
Recent advancements in CV also leverage external memory for various tasks. Several semi-parametric methods revisiting_Knn ; generalization_memory ; tip-adapter ; udandarao2022sus ; zhang2023prompt , including Tip-Adapter tip-adapter , have investigated the utilization of a $k$-nearest neighbor (kNN) classifier knn_classifier to enhance classification performance without the need for fine-tuning. RAC retrieval_long_tail integrates the output of a base model, i.e., a conventional vision encoder, with a retrieval module to address long-tail image classification challenges. Blattmann et al. retrieval_diffusion employ a semi-parametric diffusion model for generative image synthesis, which is augmented by a retrieval module.
@@
Building upon the achievements of retrieval augmentation, we introduce a framework for few-shot classification and domain generalization through retrieval-based techniques. Our work focuses on retrieving examples from few-shot training data to construct more dynamic prompts for CLIP models. The most similar method to our own is Training-free Adaption (Tip-Adapter) tip-adapter . Tip-Adapter solely employs retrieval to enhance the final classification distribution. We explore the potential of using retrieval to improve the features extracted by CLIP models.
@@
Finetuning for visual language models is crucial for bridging the domain gap in downstream tasks. While CLIP can perform zero-shot image classification by learning reasonable similarity in high-dimensional joint feature space between image-text pairs, finetuning with few-shot samples in a parameter-efficient manner remains meaningful. To achieve this, CoOP CoOP substitutes the input context token in the text branch of CLIP with learnable parameters prefix for few-shot learning. Following recent advancementsProDA ; ProGrad ; prompt_pretraining , we aim to optimize text prompts more effectively from a gradient perspective. CoCoOP CoCoOP proposes to train an intermediate network to generate image tokens as conditional inputs for the text branch and design a base-to-new benchmark to evaluate the generalization ability of a method within a dataset. Some studies CAVPT ; MaPLe ; promptsrc ; UPT ; video_prompt introduce trainable visual prompts VPT into the vision branch of CLIP to develop the dual prompt tuning scheme. Maple MaPLe proposes to learn layer-wise prompts for the two branches simultaneously. PromptSRC promptsrc utilizes its original features to regularize the learning of prompts.
@@
Adapters are typically subnetworks consisting of two fully connected layers with a nonlinear activation function in between. Referring to the concept of adaptersparameter-efficient , CLIP-Adapter clip_adapter employs lightweight adapters to adapt final features. Tip-Adapter tip-adapter utilizes a key-value cache model for auxiliary classification. Sus-X udandarao2022sus constructs a support set using generative models and extends Tip-Adapter by leveraging image-text affinity matrices. CaFO zhang2023prompt improves visual recognition in low-shot regimes by expanding the training set through a generative model and integrating prior knowledge from multiple pre-training models using multiple adapters.
@@
3 Preliminaries
CLIP CLIP is a pre-trained model that learns aligned vision-language representations from web-scale image-text datasets. It comprises two sub-networks: an image encoder $\mathrm{e}_{\mathrm{I}}$ and a text encoder $\mathrm{e}_{\mathrm{T}}$. The image encoder encodes visual inputs, and the text encoder encodes text inputs, into a shared hidden space $\mathbb{R}^{d}$. Here, $d$ represents the dimension of embeddings (e.g.,d = 512 for ViT ViT in CLIP).
@@ 3 Preliminaries
Classification of CLIP. Given an image $\boldsymbol{x}$ and a set of $C$ category names $T=\{\boldsymbol{t}_{1},\boldsymbol{t}_{2},\dots,\boldsymbol{t}_{C}\}$ (e.g., $C=1000$ for ImageNet imagenet_paper ), the image encoder extracts the image feature $\boldsymbol{z}=e_{\mathrm{I}}(\boldsymbol{x})\in\mathbb{R}^{d}$. The category names in $T$ are combined with a hand-crafted text prompt template “a photo of a [CLASS]” to generate text descriptions. These descriptions are then fed into the text encoder to derive the text features $\boldsymbol{f}\in\mathbb{R}^{d\times C}$. The prediction probability of $\boldsymbol{x}$ belonging to class $c$ is calculated using the inner product similarity between the query image feature $\boldsymbol{z}$ and text features $\boldsymbol{f}$ as shown in Equation 1:
$p(c|x)=\frac{\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)}{\sum_{c=1}^{C}\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)}.$ (1)
$\tau$ is the temperature coefficient learned by CLIP. The text encoder produces a retrieval-based dynamic classifier.
Text prompt learning. In contrast to hand-crafted prompt engineering, text prompt learning aims to generate more adaptive text features by learning a set of prompts. CoOp CoOP learns a set of parameters $\boldsymbol{P}_{\mathrm{T}}\in\mathbb{R}^{d\times M}$ to replace the predefined prompt templates, where $M$ represents the prompt length. The word token of each category name in $T$ is inserted into a template and treated as the initial values of learnable input prompts. These prompts are further fine-tuned on few-shot data to generate text features. The fine-tuned prompts $\boldsymbol{P}_{\mathrm{T}}$ adjust the decision boundaries of text features for downstream tasks. All parameters inherited from the pretrained CLIP model remain fixed during the training process.
Visual prompt learning. VPT VPT proposes a visual prompt tuning approach, which introduces a few learnable parameters into the input space while freezing the image encoder $e_{\mathrm{I}}$. This method aims to extract more transferable visual features from downstream data. In the context of an image encoder with $L$ layers, the output of the $i$-th layer, $l_{i}$, for $i=1,2,\dots,L$, can be expressed as:
@@
In this section, we propose a straightforward baseline approach for few-shot image classification, named Vision-Language Prompt Tuning (VLPT). Figure 1 illustrates the overall architecture of VLPT. The visual prompt is optimized through visual prompt learning, while the text prompt is refined via text prompt learning. Upon deriving the image and text features for each category, the model with tuned prompts generates predictions based on the classification paradigm of clip.
@@
Figure 2: The overall workflow of RePrompt includes four main steps: (a) Encoding an image input into a query embedding using a frozen image encoder; (b) Encoding each image entry from the training dataset into key and value embedding pairs with the same frozen image encoder, where the value embeddings also include one-hot representations of labels. We retrieve the top-K relevant knowledge items through maximum inner product search and integrate this knowledge to generate visual prompts; (c) Retrieval-enhanced visual prompts are introduced into the $J$ layers of the visual branch inputs, while the other prompts remain consistent with those of the baseline VLPT; (d) The final output is derived from a linear combination of the prompt-tuned CLIP prediction and the retrieval-aided prediction.
4 Proposed Method

### v3

@@ Abstract
The Contrastive Language-Image Pretraining (CLIP) model has been widely used in various downstream vision tasks. The few-shot learning paradigm has been widely adopted to augment its capacity for these tasks. However, current paradigms may struggle with fine-grained classification, such as satellite image recognition, due to widening domain gaps. To address this limitation, we propose retrieval-enhanced visual prompt learning (RePrompt), which introduces retrieval mechanisms to cache and reuse the knowledge of downstream tasks. RePrompt constructs a retrieval database from either training examples or external data if available, and uses a retrieval mechanism to enhance multiple stages of a simple prompt learning baseline, thus narrowing the domain gap. During inference, our enhanced model can reference similar samples brought by retrieval to make more accurate predictions. A detailed analysis reveals that retrieval helps to improve the distribution of late features, thus, improving generalization for downstream tasks. Reprompt attains state-of-the-art performance on a wide range of vision datasets, including 11 image datasets, 3 video datasets, 1 multi-view dataset, and 4 domain generalization benchmarks.
Index Terms:
@@
Few-shot learning, prompt tuning, image classification, video understanding, multi-view image classification
I Introduction
@@
Visual concept recognition has achieved remarkable success in closed-set scenarios with large-scale training sets, typically ImageNet dataset. The recognition accuracy can even surpass human ability. However, it may not always be feasible to have a large training set for each specific visual concept in certain downstream tasks. How to learn a robust visual concept recognition system with low-shot data has become a challenging but valuable problem in computer vision. Previous methods [1, 2, 3] mainly focus on learning a transferable visual representation from a source domain and quickly adapting to few-shot downstream tasks through fine-tuning techniques. However, limited by the categories and closed-set training samples in the source domain, these few-shot learning algorithms are only effective in impractically simple settings, such as distinguishing $5$-way $1$-shot classification.
@@
Recent advancements in visual representation learning have been propelled by the emergence of large-scale vision-language models, such as CLIP [4] and ALIGN [5]. The feasibility of applying these visual language models in addressing more challenging low-shot learning problems has attracted considerable attention within the research community [4, 6, 7, 8]. In practice, it is cumbersome to fine-tune the entire vision-language model for transfer learning in each downstream task. Furthermore, finetuning the full model can result in catastrophic forgetting [9, 10], leading to poor performance on new tasks despite the pre-trained model’s initial success. To address these challenges, CoOP [6] first proposes a few-shot evaluation framework for visual language models. There are full $C$ classes of downstream tasks and each class has 1/2/4/8/16-shot samples. Meanwhile, CoOP designs a learnable text prompt to replace the sub-optimal hand-crafted text prompt templates.
@@
Nevertheless, the frozen image feature representation also leads to sub-optimal performance. Inspired by VPT [11], many studies [12, 13, 14, 8, 15] consider adding additional learnable visual tokens into the image encoder to parameter-efficient fine-tune the vision-language models on downstream tasks. Despite significant improvements achieved by these methods for few-shot learning, these parametric models struggle to generalize to extremely low-shot data or atypical instances, often relying on rote memorization. For instance, given only one image per class, CoOP performs even worse than the zero-shot classification results of CLIP.
@@
Retrieval-augmentation approaches [16, 17, 18, 19, 20, 21] are employed to retrieve knowledge corpus and generate additional references, thereby enhancing performance in low-resource scenarios. Large Language Models (LLM) also get advancements with retrieval-augmentation approaches and further develop Retrieval Augmentation Generation (RAG) [22]. However, the application of retrieval augmentation to visual language models remains underexplored. Tip-Adapter [21], a key-value cache model, is constructed as a retrieval auxiliary classifier for CLIP. It leverages the similarity between the test and few-shot training samples, calculating affinity matrixes in text feature space and image feature space. However, it is observed that the variances and means of visual modal similarities are smaller compared to those in visual-language modalities [23]. This phenomenon is caused by CLIP maximizing the cosine similarity between paired samples in different modalities during its contrastive learning training, without considering the intra-modality similarity. Consequently, the image similarity metrics employed by Tip-Adapter, derived from CLIP, are potentially compromised by domain gaps and an insufficient adaptation to downstream datasets.
@@
In this work, we investigate the feasibility of learning prompts from a cache model, which facilitates the enhancement of query image features to adjust image-image similarities. To simultaneously optimize these aspects, we introduce Retrieval-enhanced Visual Prompt learning (RePrompt), which incorporates a retrieval mechanism for learning knowledge representations from downstream tasks. RePrompt establishes a retrieval database that stores common training image representations as retrieval keys. The current image representations retrieve relevant cached knowledge from the database. Unlike Sus-X [23], which expands the retrieval database using external datasets like LAION [24], our method adheres strictly to the principles of few-shot learning by not utilizing external datasets. Nonetheless, our method can still benefit from incorporating external knowledge.
@@ II Related Work
Retrieval-augmented models have gained significant attention in the fields of Natural Language Processing (NLP) and Computer Vision (CV). Numerous studies have augmented large language models with external memory through retrieval-enhancement mechanism [25, 19, 26, 17, 18, 27]. The integration of knowledge into a language model involves retrieving relevant samples from external memory, thereby enabling the model to generate more informed predictions based on these samples. Typically, the external memory comprises a collection of text paragraphs or a structured knowledge base. The BERT-kNN model [26] combines the outputs of a trained BERT model with a non-parametric $k$NN model for question-answering tasks. REALM [19] acquires external knowledge from Wikipedia. Karpukhin et al. [25] developed a dense passage retrieval system for open-domain question answering. RETRO [17] demonstrates how to blend retrieved tokens with input tokens in a transformer architecture and methodically evaluates the influence of large-scale external memory datasets on NLP tasks. Chen et al. [18] utilize retrieved tokens as input prompts and fine-tune the entire language model. Similar to the approach in [18], we employ retrieval samples as prompts to introduce pertinent information.
@@
Recent advancements in CV also leverage external memory for various tasks. Several semi-parametric methods [28, 29, 21, 23, 30], including Tip-Adapter [21], have investigated the utilization of a $k$-nearest neighbor (kNN) classifier [31] to enhance classification performance without the need for fine-tuning. RAC [16] integrates the output of a base model, i.e., a conventional vision encoder, with a retrieval module to address long-tail image classification challenges. Blattmann et al. [32] employ a semi-parametric diffusion model for generative image synthesis, which is augmented by a retrieval module.
@@
Building upon the achievements of retrieval augmentation, we introduce a framework for few-shot classification and domain generalization through retrieval-based techniques. Our work focuses on retrieving examples from few-shot training data to construct more dynamic prompts for CLIP models. The most similar method to our own is Training-free Adaption (Tip-Adapter) [21]. Tip-Adapter solely employs retrieval to enhance the final classification distribution. We explore the potential of using retrieval to improve the features extracted by CLIP models.
@@
Finetuning for visual language models is crucial for bridging the domain gap in downstream tasks. While CLIP can perform zero-shot image classification by learning reasonable similarity in high-dimensional joint feature space between image-text pairs, finetuning with few-shot samples in a parameter-efficient manner remains meaningful. To achieve this, CoOP [6] substitutes the input context token in the text branch of CLIP with learnable parameters [33] for few-shot learning. Following recent advancements [34, 35, 36], we aim to optimize text prompts more effectively from a gradient perspective. CoCoOP [37] proposes to train an intermediate network to generate image tokens as conditional inputs for the text branch and design a base-to-new benchmark to evaluate the generalization ability of a method within a dataset. Some studies [14, 13, 15, 12, 8] introduce trainable visual prompts[11] into the vision branch of CLIP to develop the dual prompt tuning scheme. Maple [13, 38] proposes to learn layer-wise prompts for the two branches simultaneously. PromptSRC [15] utilizes its original features to regularize the learning of prompts. Beyond their focus on few-shot classification tasks, learnable prompts exhibit remarkable adaptability across diverse downstream applications, offering a unified and efficient framework for fine-tuning foundation models to address various complex tasks [39, 40, 41, 42, 43].
@@
Adapters are typically subnetworks consisting of two fully connected layers with a nonlinear activation function in between. Referring to the concept of adapters [44], CLIP-Adapter [7] employs lightweight adapters to adapt final features. Tip-Adapter‘[21] utilizes a key-value cache model for auxiliary classification. DPT [38] propose class-aware visual prompts to guide the CLIP model focus on the target object in downstream work. Sus-X [23] constructs a support set using generative models and extends Tip-Adapter by leveraging image-text affinity matrices. CaFO [30] improves visual recognition in low-shot regimes by expanding the training set through a generative model and integrating prior knowledge from multiple pre-training models using multiple adapters.
@@
III Preliminaries
CLIP [4] is a pre-trained model that learns aligned vision-language representations from web-scale image-text datasets. It comprises two sub-networks: an image encoder $\mathrm{e}_{\mathrm{I}}$ and a text encoder $\mathrm{e}_{\mathrm{T}}$. The image encoder encodes visual inputs, and the text encoder encodes text inputs, into a shared hidden space $\mathbb{R}^{d}$. Here, $d$ represents the dimension of embeddings (e.g.,d = 512 for ViT[45] in CLIP).
@@ III Preliminaries
Classification of CLIP. Given an image $\boldsymbol{x}$ and a set of $C$ category names $T=\{\boldsymbol{t}_{1},\boldsymbol{t}_{2},\dots,\boldsymbol{t}_{C}\}$ (e.g., $C=1000$ for ImageNet [46]), the image encoder extracts the image feature $\boldsymbol{z}=e_{\mathrm{I}}(\boldsymbol{x})\in\mathbb{R}^{d}$. The category names in $T$ are combined with a hand-crafted text prompt template “a photo of a [CLASS]” to generate text descriptions. These descriptions are then fed into the text encoder to derive the text features $\boldsymbol{f}\in\mathbb{R}^{d\times C}$. The prediction probability of $\boldsymbol{x}$ belonging to class $c$ is calculated using the inner product similarity between the query image feature $\boldsymbol{z}$ and text features $\boldsymbol{f}$ as following:
$p(c|x)=\frac{\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)}{\sum_{c=1}^{C}\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)},$ (1)
where $\tau$ is the temperature coefficient learned by CLIP. The text encoder produces a retrieval-based dynamic classifier.
Text prompt learning. In contrast to hand-crafted prompt engineering, text prompt learning aims to generate more adaptive text features by learning a set of prompts. CoOp [6] learns a set of parameters $\boldsymbol{P}_{\mathrm{T}}\in\mathbb{R}^{d\times M}$ to replace the predefined prompt templates, where $M$ represents the prompt length. The word token of each category name in $T$ is inserted into a template and treated as the initial values of learnable input prompts. These prompts are further fine-tuned on few-shot data to generate text features. The fine-tuned prompts $\boldsymbol{P}_{\mathrm{T}}$ adjust the decision boundaries of text features for downstream tasks. All parameters inherited from the pretrained CLIP model remain fixed during the training process.
Visual prompt learning. VPT [11] proposes a visual prompt tuning approach, which introduces a few learnable parameters into the input space while freezing the image encoder $e_{\mathrm{I}}$. This method aims to extract more transferable visual features from downstream data. In the context of an image encoder with $L$ layers, the output of the $i$-th layer, $l_{i}$, for $i=1,2,\dots,L$, can be expressed as:
@@
In this section, we propose a straightforward baseline approach for few-shot image classification, named Vision-Language Prompt Tuning (VLPT). Fig. 1 illustrates the overall architecture of VLPT. The visual prompt is optimized through visual prompt learning, while the text prompt is refined via text prompt learning. Upon deriving the image and text features for each category, the model with tuned prompts generates predictions based on the classification paradigm of CLIP.
@@
Figure 2: The overall workflow of RePrompt includes four main steps: (a) Encoding an image input into a query embedding using a frozen image encoder; (b) Encoding each image entry from the training dataset into key and value embedding pairs with the same frozen image encoder, where the value embeddings also include one-hot representations of labels. We retrieve the top-K relevant knowledge items through maximum inner product search and integrate this knowledge to generate visual prompts; (c) Retrieval-enhanced visual prompts are introduced into the $J$ layers of the visual branch inputs, while the other prompts remain consistent with those of the baseline VLPT; (d) The final output is derived from a linear combination of the prompt-tuned CLIP prediction and the retrieval-aided prediction. The retrieval database for each dataset comes from a corresponding few-shot training set for fairness.
IV Proposed Method